40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1169     Entry time: Wed Dec 3 11:58:10 2008
Author: Alberto 
Type: Update 
Category: PSL 
Subject: MC Alignment 
Rana, Alberto,

more details on the MC alignment we did yesterday.


Last week Rana re-aligned the Mach Zender (MZ) on the PSL table to reduce the power at the dark port (see elog entry #1151). After that, the beam was aligned to the MZ but not properly aligned to the Mode Cleaner (MC) anymore. As a result the MC could not lock or did it on unwanted transverse modes. To fix that we decided to change the alignment of the MC input periscope on the PSL table.


The ultimate goal of the operation was to align the MC transmitted beam to the IFO and to maximize the power.
Such a condition depends on:
a) a good cavity alignment and
b) input beam matching to the cavity TEM00 mode.


Since the MZ alignment had only affected the input beam, we assumed the cavity alignment was still good, or at least it had not changed, and we focused on the input beam.

The IOO computer, by the MC autolocker script, is able to change the cavity alignment and the length to match the input beam and lock the cavity. Although both the length servo (LSC) and the alignment servo (WFS) have a limited effective operating range. So for the script to work properly and at best, input beam and cavity matching have to be not far from that range.

The MC periscope has two mirrors which control the pitch and yaw input angles. By changing either yaw or pitch of both mirrors together (“two-knob" technique) one can change the input angle without moving the injection point on the cavity input mirror (MC1). So this is the procedure that we followed:

  • 1) turned of the autolocker running the MC-down script
    2) brought the reflected beam spot back on the MC-reflection camera and on the reflection photodiode (REFL-PD)
    3) turned on the LSC servo
    4) tweaked the periscope's mirrors until the cavity got locked on a TEM00 mode
    5) tweaked the periscope aiming at ~0.3V from the REFL-PD and ~3V on the transmission photodiode (TRANS-PD).


Following the steps above we got ~0.5 V on the REFL-PD and ~2V on the TRANS-PD but no better than that.

Looking at the Drift Monitor MEDM screen, we found that the cavity was not in the reference optimal position, as we initially assumed, thus limiting the matching of the beam to the MC.

We restored the optics reference position and repeated the alignment procedure as above. This time we got ~3V on the TRANS-PD and ~0.5 on the REFL-PD. We thought that the reason for still such a relatively high reflection was that the beam was not well centered on the REFL-PD (high order modes pick-up?).

On the AS table we centered the REFL-PD by aligning a beam splitter in the optical path followed by the light to reach the photodiode.

We also centered the beam on the reflection Wave Front Sensors (WFS). To do that we halved the power on the MZ to reduce the sidebands power and prevent the WFS QPD from saturating. We then aligned the beam splitters on the QPD by balancing the power among the quadrants. Finally we restored the power on the MZ.

As a last thing, we also centered the transmitted beam on the TRANS-QPD.


The MC is now aligned and happily locked with 3V at the TRANS-PD and 0.3V at REFL-PD.
ELOG V3.1.3-